# Automatic Outlier Rectification via Optimal Transport

### v1

@@ 7 Conclusion
In conclusion, we propose an automatic outlier rectification mechanism that integrates outlier correction and estimation within a unified optimization framework. Our novel approach leverages the optimal transport distance with a concave cost function to construct a rectification set within the realm of probability distributions. Within this set, we identify the optimal distribution for conducting the estimation task. Notably, the concave cost function’s ”long hauls” attribute facilitates moving only a fraction of the data to distant positions while preserving the remaining dataset, enabling efficient outlier correction during the optimization process. Through comprehensive simulation and empirical analyses involving mean estimation, least absolute regression, and fitting option implied volatility surfaces, we substantiate the effectiveness and superiority of our method over conventional approaches. This demonstrates the potential of our framework to significantly enhance outlier detection integrated within the estimation process across diverse analytical scenarios.
Acknowledgments
The material in this paper is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-20-1-0397. Additional support is gratefully acknowledged from NSF 1915967, 2118199, 2229012, 2312204.
@@ Appendix A Organization of the Appendix
  • The proof of Theorem 1 and Theorem 2 are given in Section B.1.
@@
  • The long haul structure for linear regression with a concave cost function is provided in Section B.2.
  • Additional details on mean estimation simulations are given in Section C.1.
@@
  • Additional plots for the mean estimation simulations for the concave cost function are given in Section C.2.
@@
  • Additional plots for the mean estimation simulations for the convex cost function are given in Section C.3.
  • Additional details on LAD regression simulations are given in Section C.4.
@@
  • Additional plots for the LAD regression simulations for the concave cost function are given in Section C.5.
@@
  • Additional plots for the LAD regression simulations for the convex cost function are given in Section C.6.
@@
  • Additional option volatility surface experiment details are given in Section C.7.
@@
  • Additional information on the volatility surface data set and losses are given in Section C.8.
Appendix B Proof details
B.1 The proof of Theorem 1 and Theorem 2.
To start with, we give two crucial lemmas 1 and 2.
@@ Appendix B Proof details
Suppose that there is an increasing sequence $0\leq x_{1}<x_{2}<\cdots<x_{n}$. The optimal solution of
$\max_{\lambda\geq 0}\sum_{i=1}^{n}\alpha_{i}\min\{x_{i},\lambda x_{i}^{r}\}-\lambda\delta$
@@
$k:=\mathop{\max}_{k\in[n]}\left\{k:\sum_{i=k}^{n}\alpha_{i}x_{k}^{r}\geq\delta\right\}.$
@@
$x_{t}=\eta x_{t+1}<\eta\lambda x_{t+1}^{r}=\frac{\lambda}{\eta^{r}}\eta x_{t}^{r}=\lambda\eta^{1-r}x_{t}^{r}\leq\lambda x_{t}^{r}.$
@@
where $1\leq k\leq n$ and $\alpha_{i}$ are weights associated with each $x_{i}$. Two cases arise:
@@
    For all $11\leq t\leq k-$ 1, $x_{t}<\lambda x_{t}^{r}$, or for all $k+1\leq t\leq n$, $x_{t}>\lambda x_{t}^{r}$.
@@
$k=\max_{k\in[n]}\left\{\sum_{i=k}^{n}\alpha_{i}x_{i}^{r}\geq\delta\right\}.$
@@
Based on our discussion, we exclude the corner case where $k=1$. Therefore, $\delta$ is sufficiently large such that $\sum_{i=1}^{n}\alpha_{i}x_{i}^{r}\leq\delta$. It’s evident that in this trivial scenario, the optimal function value is zero.
@@
Proof of Theorem 2
B.2 Linear Regression with Concave Cost Function
In this subsection, we aim to illustrate an example as mentioned in Section 4.2. Suppose we choose a concave cost function $c$ that grows strictly slower than the loss function $f$. In such a scenario, when the budget is limited, the rectified data point $z_{\textrm{best}}$ consistently demonstrates the long-haul structure, thereby ensuring automatic outlier identification properties.
Example 2.
@@ B.2 Linear Regression with Concave Cost Function
As we shall see when the dual variable $\lambda$ is relatively small, the best rectification distribution will keep the long-haul structure although the inner minimization problem is no longer concave.
@@
$=$ $\min_{\|\Delta\|\leq\frac{|\tilde{\theta}^{T}Z^{\prime}|}{\|\tilde{\theta}\|_{*}}}\left\{(|\tilde{\theta}^{T}Z^{\prime}|-\|\tilde{\theta}\|_{*}\|\Delta\|)^{2}+\lambda\|\Delta\|^{r}\right\}.$
@@
    When $\lambda\geq 4\left(|\tilde{\theta}^{T}Z^{\prime}|\|\tilde{\theta}\|_{*}\sqrt{\frac{|\tilde{\theta}^{T}Z^{\prime}|}{3\|\tilde{\theta}\|_{*}}}-\left(\sqrt{\frac{|\tilde{\theta}^{T}Z^{\prime}|}{3\|\tilde{\theta}\|_{*}}}\right)^{3}\right)$ (i.e., $\delta$ is sufficiently small), we have $g(\beta^{\star})\geq 0$. As such, the critical point of the unconstrained optimization problem is not in the interval $[0,\frac{|\tilde{\theta}^{T}Z^{\prime}|}{\|\tilde{\theta}\|_{*}}]$. In other words, we can conclude the optimal solution will be 00 and the solution of vanilla least square is already optimal.
@@
Appendix C Additional Experiments & Experimental Details
C.1 Mean Estimation Details

### v2

@@ 6 Conclusion
In conclusion, we propose an automatic outlier rectification mechanism that integrates outlier correction and estimation within a unified optimization framework. Our novel approach leverages the optimal transport distance with a concave cost function to construct a rectification set within the realm of probability distributions. Within this set, we identify the optimal distribution for conducting the estimation task. Notably, the concave cost function’s "long hauls" attribute facilitates moving only a fraction of the data to distant positions while preserving the remaining dataset, enabling efficient outlier correction during the optimization process. Through comprehensive simulation and empirical analyses involving mean estimation, least absolute regression, and fitting option implied volatility surfaces, we substantiate the effectiveness and superiority of our method over conventional approaches. This demonstrates the potential of our framework to significantly enhance outlier detection integrated within the estimation process across diverse analytical scenarios.
@@ Appendix A Organization of the Appendix
  • The organization of the Appendix is given in this section (Section A).
  • Additional details on the connection between our estimator and prior work are given in Section B.1.
  • Additional explanation of the long haul structure induced by our estimator and extensions to more general forms of problems are given in Appendix B.2.
  • The proof of Theorem 2 and Theorem 4 are given in Section C.1.
@@
  • The long haul structure for linear regression with a concave cost function is provided in Section C.2.
  • The detailed computational procedure for mean estimation, linear absolute regression, and more general cases can be found in Section D.
  • The main experiments for mean estimation and LAD regression are described in Appendix E.1 and E.2 respectively. Additional details are given in the following appendices.
  • Additional details on the illustrative example in the text is given in Section F.1.
  • Additional details on mean estimation simulations are given in Section F.2.
@@
  • Additional plots for the mean estimation simulations for the concave cost function are given in Section F.3.
@@
  • Additional plots for the mean estimation simulations for the convex cost function are given in Section F.4.
  • Additional details on LAD regression simulations are given in Section F.5.
@@
  • Additional plots for the LAD regression simulations for the concave cost function are given in Section F.6.
@@
  • Additional plots for the LAD regression simulations for the convex cost function are given in Section F.7.
@@
  • Additional option volatility surface experiment details are given in Section G.1.
@@
  • Additional information on the volatility surface data set and losses are given in Section G.2.
  • Background and additional details on the deep learning volatility modeling experiment are given in Section G.3.
Appendix B Estimator Remarks
B.1 Connections to Prior Work
Additional details on connections to prior work are given in the following remark:
Remark B.1.
(i) When we disregard the outer minimization concerning estimation parameters, the inner minimization problem over probability measures is related to the minimum distance functionals-based estimator, initially proposed by [Donoho and Liu, 1988a]. The estimator is obtained by first projecting the corrupted distribution $\mathbb{P}^{\prime}_{n}$ onto a family of distributions $\mathcal{G}$ under a certain distribution discrepancy measure $\mathds{D}$. Then, the optimal parameters are selected for the resulting distribution. However, even with additional information about our contamination models, this “two-stage” procedure has two major drawbacks: the difficulty in choosing an appropriate family of distributions $\mathcal{G}$ and the inherent computational challenge to project probability distributions. Moreover, the first stage (projection) is usually sensitive to the choice of family $\mathcal{G}$. In contrast, we propose a novel approach that integrates outlier rectification (i.e., explicit projection) and parameter estimation into a joint optimization framework in (3). (ii) This min-min strategy has also been explored in Jiang and Xie [2024] using an artificially constructed rectification set. Their primary focus is to utilize the min-min formulation to recover well-studied robust statistics estimators. In contrast, our focus is on a new conceptual framework for formulating novel estimators which have not been studied before.
B.2 Long Haul Structure and Extensions
An explanation of the long haul structure and extensions to more general forms are given in this subsection.
(i) (Concave cost function) From Proposition 1, the optimal rectification admits $z_{\textrm{best}}\in\arg\min_{z\in\mathcal{Z}}\ell(\theta,z^{\prime})+\lambda c(z,z^{\prime}).$ Suppose we select a concave cost function $c$ that grows strictly slower than the loss function $\ell$. In this scenario, when the budget is small, the rectified data point $z_{\textrm{best}}$ consistently exhibits the long haul structure, ensuring automatic outlier identification properties. This implies that the rectification process effectively identifies outliers, as the influence of the cost function dominates over the loss function for small budgets. To further illustrate the concept, we can consider a linear regression problem with the squared loss function $\ell(\theta,z)=(y-\theta^{T}x)^{2}$ or a nonlinear loss using the $r$-th norm as the cost function. We refer the readers to Appendix C.2.
(ii) (Other distribution metric/discrepancy) Based on the two examples discussed earlier, we can infer that the optimal transport distance-based rectification set is particularly suitable for regression or problems with continuous response variables. This is because the budget $\delta$ is used to compensate for the loss caused by identified outliers. However, in classification tasks, where the loss function may be less sensitive, the effectiveness of the optimal transport distance-based approach may be limited. To address this limitation and complement the rectification set for classification tasks, an alternative approach is to use $\phi$-divergences. This approach has already been proposed in a study by [Chen and Royset, 2022, Antil et al., 2024] for handling outliers in image classification tasks. The $\phi$-divergence-based rectification set operates by adjusting the weight assigned to outliers, effectively minimizing their impact on the overall data set. By reducing the weight placed on the detected outliers, the rectification set aims to remove their influence from the data set. This selective adjustment allows for the identification and removal of outliers, leading to a more refined and reliable data set.
Appendix C Proof Details
C.1 Proof of Theorem 2 and Theorem 4.
To start with, we give two crucial lemmas 5 and 6.
@@ Appendix C Proof Details
Suppose that there is an increasing sequence $0\leqslant x_{1}<x_{2}<\dots<x_{n}$. The optimal solution of
$\max_{\lambda\geqslant 0}\sum_{i=1}^{n}\alpha_{i}\min\{x_{i},\lambda x_{i}^{r}\}-\lambda\delta$
@@
$k:=\mathop{\max}_{k\in[n]}\left\{k:\sum_{i=k}^{n}\alpha_{i}x_{k}^{r}\geqslant\delta\right\}.$
@@
$x_{t}=\eta x_{t+1}<\eta\lambda x_{t+1}^{r}=\frac{\lambda}{\eta^{r}}\eta x_{t}^{r}=\lambda\eta^{1-r}x_{t}^{r}\leqslant\lambda x_{t}^{r}.$
@@
where $1\leqslant k\leqslant n$ and $\alpha_{i}$ are weights associated with each $x_{i}$. Two cases arise:
@@
    For all $11\leqslant t\leqslant k-$ 1, $x_{t}<\lambda x_{t}^{r}$, or for all $k+1\leqslant t\leqslant n$, $x_{t}>\lambda x_{t}^{r}$.
@@
$k=\max_{k\in[n]}\left\{\sum_{i=k}^{n}\alpha_{i}x_{i}^{r}\geqslant\delta\right\}.$
@@
Based on our discussion, we exclude the corner case where $k=1$. Therefore, $\delta$ is sufficiently large such that $\sum_{i=1}^{n}\alpha_{i}x_{i}^{r}\leqslant\delta$. It is evident that in this trivial scenario, the optimal function value is zero.
@@
We now give the proof of Theorem 2:
Proof.
Before we prove the theorem, it is worth highlighting that for any fixed $\theta$, we can always sort $\{z_{i}\}_{i=1}^{n}$ based on the error $\|\theta-z_{i}^{\prime}\|$ to satisfy condition (6).
By the strong duality result in Proposition 1, we have
=${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}{=}}\min\limits_{\mathbb{Q}\in\mathcal{B}(\mathbb{P}^{\prime}_{n})}\mathbb{E}_{\mathbb{Q}}[\|\theta-Z^{\prime}\|]$
$=\max_{\lambda\geqslant 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}\left[\min_{\Delta\in\mathbb{R}}\|\theta-Z^{\prime}-\Delta\|+\lambda\|\Delta\|^{r}\right]-\lambda\delta$
$=\max_{\lambda\geqslant 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}\left[\min_{0\leqslant\|\Delta\|\leqslant\|\theta-Z^{\prime}\|}\|\theta-Z^{\prime}\|-\|\Delta\|+\lambda\|\Delta\|^{r}\right]-\lambda\delta$
$=\max_{\lambda\geqslant 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}[\min\left\{\|\theta-Z^{\prime}\|,\lambda\|\theta-Z^{\prime}\|^{r}\right\}]-\lambda\delta$
$=\max_{\lambda\geqslant 0}\frac{1}{n}\sum_{i=1}^{n}\min\{\|\theta-z_{i}^{\prime}\|,\lambda\|\theta-z_{i}^{\prime}\|^{r}\}-\lambda\delta$
where the third equality follows from Lemma 5 in Appendix C.1 and the fifth one is due to Lemma 6 in Appendix C.1. ∎
The proof in the LAD regression case follows:
Proof.
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}=}\min\limits_{\mathbb{Q}\in\mathcal{R}(\mathbb{P}^{\prime}_{n})}\mathbb{E}_{\mathbb{Q}}[\|Y-\theta^{T}X\|]$
$=\max_{\lambda\geqslant 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}\left[\min_{\Delta\in\mathbb{R}^{d+1}}\|\tilde{\theta}^{T}Z^{\prime}+\tilde{\theta}^{T}\Delta\|+\lambda\|\Delta\|^{r}\right]-\lambda\delta$
$\mathop{=}\limits^{(a)}\max_{\lambda\geqslant 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}\left[\min_{\|\Delta\|\leqslant\frac{\|\tilde{\theta}^{T}Z^{\prime}\|}{\|\tilde{\theta}\|_{*}}}\|\tilde{\theta}^{T}Z^{\prime}\|-\|\tilde{\theta}\|_{*}\|\Delta\|+\lambda\|\Delta\|^{r}\right]-\lambda\delta$
$\mathop{=}\limits^{(b)}\max_{\lambda\geqslant 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}\left[\min\left\{\|\tilde{\theta}^{T}Z^{\prime}\|,\frac{\lambda\|\tilde{\theta}^{T}Z^{\prime}\|^{r}}{\|\tilde{\theta}\|_{*}^{r}}\right\}\right]-\lambda\delta$
$=\max_{\lambda\geqslant 0}\frac{1}{n}\sum_{i=1}^{n}\min\left\{\|\tilde{\theta}^{T}z_{i}^{\prime}\|,\frac{\lambda\|\tilde{\theta}^{T}z_{i}^{\prime}\|^{r}}{\|\tilde{\theta}\|_{*}^{r}}\right\}-\lambda\delta$
[…]
This completes our proof.
C.2 Linear Regression with Concave Cost Function
In this subsection, we aim to illustrate an example as mentioned in Section 4.2. Suppose we choose a concave cost function $c$ that grows strictly slower than the loss function $f$. In such a scenario, when the budget is limited, the rectified data point $z_{\textrm{best}}$ consistently demonstrates the long haul structure, thereby ensuring automatic outlier identification properties.
Example C.1.
@@ C.2 Linear Regression with Concave Cost Function
As we shall see when the dual variable $\lambda$ is relatively small, the best rectification distribution will keep the long haul structure although the inner minimization problem is no longer concave.
@@
$=$ $\min_{\|\Delta\|\leqslant\frac{|\tilde{\theta}^{T}Z^{\prime}|}{\|\tilde{\theta}\|_{*}}}\left\{(|\tilde{\theta}^{T}Z^{\prime}|-\|\tilde{\theta}\|_{*}\|\Delta\|)^{2}+\lambda\|\Delta\|^{r}\right\}.$
@@
    When $\lambda\geqslant 4\left(|\tilde{\theta}^{T}Z^{\prime}|\|\tilde{\theta}\|_{*}\sqrt{\frac{|\tilde{\theta}^{T}Z^{\prime}|}{3\|\tilde{\theta}\|_{*}}}-\left(\sqrt{\frac{|\tilde{\theta}^{T}Z^{\prime}|}{3\|\tilde{\theta}\|_{*}}}\right)^{3}\right)$ (i.e., $\delta$ is sufficiently small), we have $g(\beta^{\star})\geqslant 0$. As such, the critical point of the unconstrained optimization problem is not in the interval $[0,\frac{|\tilde{\theta}^{T}Z^{\prime}|}{\|\tilde{\theta}\|_{*}}]$. In other words, we can conclude the optimal solution will be 00 and the solution of vanilla least square is already optimal.
@@
Appendix D Computational Procedure
D.1 Alternative Estimation Procedure for Inner Problem
We make the following remark on an alternative procedure for estimating $k(\theta_{t})$ and $\lambda$ in the inner problem for mean estimation and LAD regression in Algorithm 1:
Remark D.1 (Alternative procedure for $k(\theta_{t})$ and $\lambda^{*}$).
It is illuminating to note that steps 1 and 2 in Algorithm 1 can be alternatively replaced with steps that solve a specific linear program for the knot point $k(\theta_{t})$ and optimal $\lambda^{*}$. Instead of applying the sorting and quick-select algorithm implied by Lemma 6, we can instead reformulate Equation (8) as a linear programming problem, i.e.,
$\begin{array}[]{ccll}&\max\limits_{\lambda\geqslant 0,t\in\mathbb{R}^{n}}&\frac{1}{n}\sum_{i=1}^{n}t_{i}-\lambda\delta\\ &\mathrm{s.t.}&t_{i}\leqslant\|\tilde{\theta}^{T}z_{i}^{\prime}\|,\forall i\in[n]\\ &&\|\tilde{\theta}\|_{*}^{r}t_{i}\leqslant\lambda\|\tilde{\theta}^{T}z_{i}^{\prime}\|^{r},\forall i\in[n].\end{array}$ (9)
The problem in (9) can then be solved instead of performing steps 1 and 2 of Algorithm 1 to find the knot point $k(\theta_{t})$ and the optimal $\lambda^{*}$.
D.2 General Computational Procedure
We also propose a procedure for fitting more general regression models with our statistically robust estimator, which is based on the approach for LAD regression above.
$\max_{\lambda\geqslant 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}\left[\min_{z\in\mathcal{Z}}\ell(\theta;z)+\lambda c(z,Z^{\prime})\right]-\lambda\delta.$ (10)
Although we do not prove a reformulation result for this problem, we solve the problem empirically by employing a computational procedure which is analogous to that of LAD regression. The procedure is heuristic and does not enjoy the same guarantees as Algorithm 1 for LAD regression, but in experiments (see Section 5.2.2), we find that the performance is similarly robust to outliers and significantly outperforms benchmark models trained under empirical risk minimization. Our computational procedure follows in Algorithm 2.
Data: Observed data $\{{z_{i}}^{\prime}\}_{i=1}^{n}$, initial point $\theta^{(0)}$, stepsizes $\alpha^{(t)}>0$, sampling distribution $\mathbb{P}_{n}^{\prime}$; batch size $m\leqslant n$.
1 for $t=0,\dots,T$ do
2       1. Sample $m$ points $\{z_{i}^{\prime}\}_{i=1}^{m}\sim\mathbb{P}_{n}^{\prime}$.
5       4. Subgradient step on the detected clean data:
6 end for
Algorithm 2 Statistically Robust Optimization Procedure
Algorithm 2 is similar to Algorithm 1, but has two principal differences: (1) instead of optimizing over the entire data set, we optimize over mini-batches of size $m\leqslant n$ which are drawn from a data set of size $n$ via a sampling distribution $\mathbb{P}_{n}^{\prime}$; and (2) instead of taking the subgradient with respect to $\theta$ for the LAD regression problem in calculating the direction of descent, we instead take a local subgradient with respect to $\theta$ around each point $z^{\prime}_{i}$, each of which is contained within the subdifferetial $\partial_{\theta}\ell(\theta^{(t)},z^{\prime}_{i})$. The subgradient is in practice computed by the automatic differentiation capability of a software package such as Pytorch, Tensorflow, or JAX. The budget parameter $\delta$, which controls the quantile which identifies outliers, is tuned via cross-validation.
D.3 Optimization Details
This issue becomes even more critical when dealing with our estimation problem, as these problems may lack weak convexity or subdifferential regularity properties. One significant challenge in such problems is the non-coincidence of different subdifferential concepts, such as Clarke, Fréchet, and limiting subdifferentials. Moreover, some calculus rules (summation and chain rules) do not always hold for these concepts. Thus, computing a first-order oracle, which provides necessary information for optimization algorithms, can be difficult in practice.
To concretely illustrate this challenge, we note that, intuitively, the graph of a subdifferentially regular function cannot exhibit "downward-facing cusps" [Li et al., 2020, Davis et al., 2020]. Here, we give an example of when the objective function (7) is not regular, which may occur in general.
Example D.1 (Irregular Case (Three Point Masses)).
Figure 5: Irregular objective function.
From Figure 5, it is evident that there are at least two regions with "downward-facing cusps" (marked as irregularities in the figures), which could potentially pose computational challenges. Similar challenges arise in the training of deep neural networks, as demonstrated in Figure 4 of [Li et al., 2020], particularly for two-layer neural networks. However, the "subgradient" method remains effective for addressing these challenges empirically. In our empirical investigation, we have observed a similar phenomenon for our estimator.
Appendix E Mean Estimation and LAD Regression Experiments
In this section, we outline the main results of the mean estimation and LAD regression experiments. Additional details are deferred to the following appendix section (Appendix F) for the convenience of readers.
E.1 Mean Estimation
Table 3: We compared our estimator with several standard mean estimation methods by evaluating the average loss on clean data points across various corruption levels. In our evaluation, we set the percent level of trimmed mean equal to the unknown ground truth corruption level. The hyperparameters for our estimator, namely $\delta=0.5$ and $r=0.5$, remained constant across all corrupted levels. The last row in the comparison table represents the percentage of all points rectified by our method. Error bars represent two standard deviation confidence intervals we have computed manually assuming normal errors.
Corruption Levels 20% 30% 40% 45% 49%
[…]
Ours, %Rectified 10.03% 10.12% 10.24% 10.35% 10.55%
We visualize the corrupted distribution and its rectified distribution when the corruption level is 45% in Figure 6. In this case, 55% of the data is drawn from the true distribution $\mathcal{N}(0,2)$ and 45% is drawn from $\mathcal{N}(25,2)$. Figure 6(a) displays the original sample of data from the contaminated distribution. The outlier points from the contaminating distribution are clearly visible in orange. Figure 6(b) shows the rectified distribution our estimator produces for $\delta=2.5$, in which the outlier points have been moved from their original values to their rectified values, which is much closer to the true mean of the clean distribution. Our estimator thus successfully identifies the majority of outliers and relocates them towards the center (our mean estimator), providing further support for our theoretical findings in the previous section. The sensitivity analysis of mean estimation with respect to $\delta$ is displayed in Figure 6(c). In this figure, the loss on the clean data is plotted for various values of $\delta$. We see that an approximate minimum occurs at $\delta=2$, with good performance within the range $\delta\in[0.5,2.5]$. This illustrates the relative insensitivity of our estimator to different values of $\delta$ for a wide and reasonable range. This range can be easily reached by hyperparameter tuning via cross-validation, as the function describing the performance of our estimator is approximately quasiconvex. Moreover, when $\delta=0$ or when $\delta$ is set so large that it can rectify all points, our estimator gracefully degrades to the loss of the median estimator, which is another favorable property. A further sensitivity analysis with respect to $r$ is given in Appendix F.2. In this sensitivity analysis we find a similarly wide region of good performance for $r$ which improves significantly and almost uniformly on the more typical setting of $r=1$.
The evolution of the rectified distribution for mean estimation under concave and convex cost functions as $\delta$ increases is depicted in Appendices F.3 and F.4. These depictions show that our statistically robust estimator which applies the concave transport cost function acts in a stable and expected manner as $\delta$ increases. In particular, this depiction shows that our estimator moves points in the order of their distance from the estimate of the mean; that is, the budget is “used” on the worst outliers first. This results in an orderly procedure of rectification which is stable and predictable across nearby values of $\delta$, which is a favorable property if $\delta$ is being set sequentially or in cross-validation. In contrast, the estimator which applies the convex transport cost function moves all points of the clean and outlier distributions toward each other, which results in poor estimates of the true mean and an improper rectified distribution.
E.2 LAD Regression
As we discussed in the previous result, the theoretical analysis for the LAD estimator resembles that of the mean estimation task. Empirically, we also observe similar performance for the LAD estimator and show that our estimator correctly rectifies most of outliers to the fitted hyperplane under a range of choices for $\delta$. In order to further support the effectiveness of our estimator, we provide additional visualizations: Figure 7 contains a visualization of the lines of best fit produced by various estimations to the LAD regression problem, and displays the effect of different choices of $\delta$ on the line of best fit produced by our estimator. In Figure 7(a) our estimator with $\delta=1$ successfully produces a line of best fit which is closest to the uncontaminated distribution, while the other estimators (OLS, LAD, and Huber regressors) produce lines which are heavily affected by the contaminating distribution, showcasing the robustness of our estimator. Figure 7(b) and 7(c) show the rectified distribution produced by our estimator under $\delta=1$ (b) and $\delta=1.5$ (c). The suboptimal value of $\delta=1$ still rectifies many points from the contaminated distribution and produces a good line of best fit. Setting $\delta=1.5$ rectifies all of the points from the contaminated distribution and essentially recovers the true line of best fit. Importantly, this simulation shows that even improperly setting $\delta$ to the suboptimal value of $\delta=1$ produces a much better line than any of the other estimators in Figure 7(a). Appendix F.5 contains comparisons over different random trials and a sensitivity analysis with respect to $\delta$ and $r$, along with experimental results.
Appendix F Additional Mean Estimation and LAD Regression Experimental Details
F.1 Illustrative Example Details
In this example, we perform linear regression using our estimator with either a concave or a convex transport cost function. The estimation is performed on a sample of points in $\mathbb{R}^{2}$ drawn from a contaminated distribution (corresponding to $\mathbb{P}^{\prime}$) where 45% of the points are distributed according to an outlier distribution and 55% of the points are distributed according to the true uncontaminated distribution (corresponding to $\mathbb{P}_{\star}$). Our goal is to recover the line of best fit which corresponds to the uncontaminated distribution, which is $y=0.1-0.8x$. In other words, 55% of the points are drawn with $x\sim\mathcal{N}(0,2)$ and $y=0.1-0.8x+0.2w$, and 45% of the points are drawn with $x\sim\mathcal{N}(4,2)$ and $y=10.1-0.8x+0.2w$. We illustrate the considerable differences which can occur when using each of these cost functions under various choices of $\delta$, the rectification budget parameter. We depict this example in Figure 2 for the concave cost and Figure 3 for the convex cost.
The behavior of the esstimator using the concave cost function is depicted in Figures 2(a)-(c). Each of these subfigures shows the points drawn from each distribution, the rectified distribution, the true line of best fit for the uncontaminated distribution, and the line of best fit produced by our estimator, at various choices of $\delta$. In Figure 2(a), $\delta=0$, in Figure 2(b), $\delta=0.9$, and in Figure 2(c), $\delta=1.5$. As can be seen from the figures, as the budget $\delta$ increases, we note that the proposed estimator using the concave cost function becomes increasingly adept at mitigating the influence of outlier data. It achieves this by rectifying the outlying points, moving them into the bulk of the uncontaminated distribution. As a result, the line of best fit (depicted in red) aligns much more closely with the true line of best fit (depicted in black). It is able to perform this successful estimation because each outlier point can be moved far with a cheap cost (a “long haul”) due to the use of the concave transport cost function.
However, the convex cost function may result in suboptimal estimators, which can only move every data point (clean and contaminated) a little bit. Figures 3(a)-(c) display the behavior of the estimator using the convex cost function, which illustrates this effect. Each subfigure displays the rectified distribution as yellow-green points for a different setting of $\delta$ (i.e. (a) $\delta=0$, (b) $\delta=0.9$, and (c) $\delta=1.5$). As can be seen from the evolution of the rectified distribution, as $\delta$ increases, instead of the outlier points being moved towards the bulk of the distribution, all of the points move towards each other–even the points in the clean distribution. This is not ideal behavior, as the points from the clean distribution should stay in place. This defective behavior causes the line of best fit produced by the estimator using the convex cost function (which is depicted in red) fit to the rectified points to be severely affected by the outliers; instead of being close to the ground truth line of best fit (which is depicted in black), its slope is moved significantly upwards towards the outliers. This occurs because the convex cost function gives lower cost to smaller rectifications of the given data set, which is an inappropriate assumption for data sets containing outliers.
Appendices F.6 and F.7 depict the evolution of the rectified distribution as $\delta$ increases for the concave and convex cost functions, respectively. For our estimator which uses the concave cost, the evolution of the rectified distribution as $\delta$ increases shows that our estimator increasingly rectifies the points in order of their distance from the line of best fit in a predictable and stable way, which is a favorable property for setting $\delta$ via hyperparameter tuning. In contrast, the convex estimator consistently moves all points (clean and corrupted) toward each other as $\delta$ changes, which achieves a poor estimate of the true line of best fit for all $\delta$ considered.
F.2 Mean Estimation Details